# Measuring the (Un)Faithfulness of Concept-Based Explanations

### v3

@@ 6.2 Benchmarking U-CBEMs
We evaluate six prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, MCD, HU-MCD). We benchmark U-CBEMs on a varied range of tasks and models to demonstrate the applicability of SURF. Specifically, our tasks are: 1) Object Classification, 2) Multi-Attribute Prediction, and 3) Age Regression. In all tasks, U-CBEMs discover 5 CAVs (or subspaces for MCD and HU-MCD) per output. CAVs and importances are found on the training set; the resulting explanations are evaluated on the test set. Dataset, finetuning, U-CBEM implementation details, and good-faith efforts made to adapt U-CBEMs to new tasks are in Sec. B.
@@
We evaluate on three tasks. (1) Object Classification: We use the Caltech-101-finetuned ResNet-50 and report $\text{SURF}_{\text{MAE}}$, $\text{SURF}_{\text{EMD}}$, Top-1, and Rank Corr. HU-MCD requires architectural modifications, so we can only evaluate it on this task, which uses a compatible ResNet-50. (2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [Sandler2018MobileNetV2IR] to predict the presence/absence of attributes on the CelebA dataset [celeba]. We report $\text{SURF}_{\text{MAE}}$ and attribute prediction accuracy (Attr-Acc), which is introduced as a task-specific replacement for Top-1. (3) Age Estimation: Given a human face, we finetune a ViT [VIT] to predict their age (in years) on the UTK-Face dataset [utkface]. Because ViTs have negative activations, U-CBEMs with non-negativity assumptions (i.e., ICE and CRAFT) cannot be used. C-SHAP is omitted as it is only applicable to classification, and MCD is incompatible with ViTs.
@@
Benchmark results are reported in Tab. 4. In the evaluated settings, we observe that no prior U-CBEM is faithful to the original model. In the Object Classification task, the most faithful U-CBEM is HU-MCD with an $\text{SURF}_{\text{EMD}}$ of 0.384, denoting large errors in the probability-space. Only two of the tested U-CBEMs (ICE and MCD) perform significantly above random chance in the Multi-Attribute Prediction task. CDISCO exhibits an average error of 32.6 years in the Age Estimation task.
@@
We believe that U-CBEMs are unfaithful for two reasons: (1) Other than C-SHAP, all methods discover class-specific CAVs to reconstruct the embedding. Crucially, the class-specific CAVs are found only on class-specific images, so using class $c_{1}$’s CAVs to reconstruct an embedding from an image of class $c_{2}$ will have large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large logit- and probability-space errors. This issue is exacerbated for CRAFT and ICE, which non-linearly project to the concept space. Though MCD and HU-MCD also find class-specific concepts, they achieve superior faithfulness by using concept subspaces; thus they faithfully explain more of the model with the same number of concepts. (2) Most methods do not satisfy a logit-based completeness criterion, preventing a simple reproduction of model outputs. Note that C-SHAP satisfies an accuracy-based completeness criterion and ICE, MCD, and HU-MCD satisfy the logit-based criterion by including uninterpretable residuals, which cannot be presented to the user in the explanation.
@@
The most important hyperparameter for any U-CBEM is the number of concepts to be discovered. Intuitively, having more concepts should result in a more complete (i.e., faithful) explanation. However, explanations containing too many concepts may be difficult to interpret. Cognitive psychology studies show that humans can hold a limited number of items in their working memory at a time [2010TheMM, Miller1956TheMN]. Thus, simple, or parsimonious, explanations are preferred, which communicate the bulk of the model’s computation with a few concepts. However, inadequate faithfulness measures make the faithfulness-parsimony tradeoff hard to analyze, so prior U-CBEMs commonly set this hyperparameter arbitrarily (10 in ICE, 25 in CRAFT & ACE). SURF allows us to intelligently analyze this tradeoff.
Using the Object Classification task from Sec. 6.2, we fit U-CBEMs with an increasing number of concepts. Each U-CBEM is evaluated using the SURF measure, visualized in Fig. 2. We observe that some U-CBEMs (CDISCO, CRAFT, C-SHAP) either do not improve or perform worse on one or both metrics as they discover more concepts. ICE marginally improves as we add more concepts. Only MCD and HU-MCD uniformly improve as the number of concepts increases, and they exhibit a plateauing effect, marking a natural choice for the number of concepts. Interestingly, HU-MCD is more faithful than MCD only when both discover relatively few concepts; as the number of concepts grows, MCD becomes more faithful. This offers a more nuanced perspective on the findings from the HU-MCD paper, where they claimed superior faithfulness over MCD based on results from their deletion-based proxy.
@@ 7 Conclusion
This paper argues that we lack a clear view on how faithful current U-CBEMs really are, largely because of measures that deviate from the formal definition of faithfulness. In accordance with this definition, we develop desiderata for a faithfulness measure and propose SURF as a simple, principled measure for faithfulness. SURF passes a basic sanity check, reveals that state-of-the-art U-CBEMs are substantially unfaithful, and is a practical tool for studying the tradeoff between parsimony and faithfulness when choosing the number of concepts. We urge future work on concept-based explanations to adopt SURF as a standard faithfulness measure and to report SURF scores alongside their interpretability claims.
[…]
@@
    Sec. A: Implementation details for the measure-over-measure study (Sec. 6.1).
@@
    Sec. B: Experimental details for the evaluation of U-CBEMs done in Sec. 6.2.
A Implementation Details for Measure-over-Measure Comparison
@@
A.1 Experimental Setting Details
@@ A.2 Surrogate Implementation
A.2.1 SURF Surrogate (Ours)
@@ A.2.3 C-SHAP Surrogate
Since C-SHAP’s surrogate contains learnable parameters. Following [ConceptSHAP], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities. We train the surrogate using the Adam optimizer [kingma2014adam] and a starting learning rate of $0.1$. The surrogate is trained for at most 100 epochs, during which the learning rate is decayed by a factor of 0.5 when the training loss plateaus. We stop training when the learning rate is reduced below $10^{-7}$.
A.3 FLOPs Computation
@@
B Additional Details for Benchmarking U-CBEM Faithfulness
@@ B.4 CDISCO
B.5 ICE
The open-source implementation provided by the authors was used to evaluate this approach. ICE uses NMF to find CAVs. Due to the potentially large number of embeddings, we use the batched-version of NMF. Concept importance is found using the TCAV method proposed in [kimTCAV]. Embeddings are projected to the concept space using sci-kit learn’s transform method. ICE was extended to multi-attribute prediction by reformulating the task as independent binary classification tasks. Thus, CAVs and concept importances were learned for each task independently.
B.6 CRAFT
The open-source implementation provided in the Xplique toolbox [fel2022xplique] (by the same authors) was used to evaluate this approach. CRAFT uses NMF to find CAVs and introduces a Sobol-based sensitivity approach to find concept importance. Embeddings are projected to the concept space using the toolbox’s built-in transform method.
B.7 ConceptSHAP
@@
    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([shapley]) and is class-specific.
@@
B.8 MCD
Instead of representing concepts as single CAVs, MCD represents concepts as multi-dimensional subspaces. In [vielhaben2023multidimensional], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace. To ensure the union of all concepts span the entire feature space, they define the orthogonal complement concept $C^{i,\perp}=\text{span}(C^{i,1},...,C^{i,K})^{\perp}$, where $K$ is the number of concepts originally discovered, letting $C^{i,K+1}\equiv C^{i,\perp}$. Then, any embedding $\mathbf{h}$ can be uniquely decomposed onto the concept basis as:

### v4

@@ 6.2 Benchmarking U-CBEMs
We evaluate seven prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, MCD, HU-MCD, SAE) on a varied range of tasks and models to demonstrate the applicability of SURF. Specifically, our tasks are: 1) Object Classification, 2) Multi-Attribute Prediction, and 3) Age Regression. In all tasks, U-CBEMs discover 5 CAVs (or subspaces for MCD and HU-MCD) per output. CAVs and importances are found on the training set; the resulting explanations are evaluated on the test set. Dataset, finetuning, U-CBEM implementation details, and good-faith efforts made to adapt U-CBEMs to new tasks are in Appendix Sec. C.
@@
We evaluate on three tasks. (1) Object Classification: We use the Caltech-101-finetuned ResNet-50 and report $\text{SURF}_{\text{MAE}}$, $\text{SURF}_{\text{EMD}}$, Top-1, and Rank Corr. HU-MCD requires architectural modifications, so we can only evaluate it on this task, which uses a compatible ResNet-50. Results on a VGG and InceptionV3 are in Appendix Sec. D. (2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [40] to predict the presence/absence of attributes on the CelebA dataset [32]. We report $\text{SURF}_{\text{MAE}}$ and attribute prediction accuracy (Attr-Acc), which is introduced as a task-specific replacement for Top-1. (3) Age Estimation: Given a human face, we finetune a ViT [9] to predict their age (in years) on the UTK-Face dataset [54]. Because ViTs have negative activations, U-CBEMs with non-negativity assumptions (i.e., ICE and CRAFT) cannot be used. C-SHAP is omitted as it is only applicable to classification, and MCD is incompatible with ViTs.
@@
Benchmark results are reported in Tab. 4. In the evaluated settings, we observe that no prior U-CBEM is faithful to the original model. In the Object Classification task, the most faithful U-CBEM is SAE with an $\text{SURF}_{\text{EMD}}$ of 0.195, denoting significant errors in the probability-space. Only half of the tested U-CBEMs (ICE, MCD, and SAE) perform significantly above random chance in the Multi-Attribute Prediction task. SAE exhibits an average error of 3.67 years in the Age Estimation task. Note that this evaluation solely focuses on faithfulness and does not make any judgment on the interpretability side of the tradeoff.
@@
We believe that U-CBEMs are unfaithful for two reasons: (1) Other than C-SHAP and SAE, all methods discover class-specific CAVs to reconstruct the embedding. Crucially, the class-specific CAVs are found only on class-specific images, so using class $c_{1}$’s CAVs to reconstruct an image embedding of class $c_{2}$ will have large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large SURF errors. This issue is exacerbated for CRAFT and ICE, which non-linearly project to the concept space. Though MCD and HU-MCD also find class-specific concepts, they achieve superior faithfulness by using concept subspaces; thus they faithfully explain more of the model with the same number of concepts. (2) Most methods do not satisfy a logit-based completeness criterion, preventing a targeted reproduction of model outputs. C-SHAP, instead, satisfies an accuracy-based completeness criterion, and ICE, MCD, and HU-MCD satisfy the logit-based criterion by including uninterpretable residuals, which cannot be shown in the user explanation.
@@
The most important hyperparameter for any U-CBEM is the number of concepts to be discovered. Intuitively, having more concepts should result in a more complete (i.e., faithful) explanation. However, explanations containing too many concepts may be difficult to interpret. Cognitive psychology studies show that humans can hold a limited number of items in their working memory at a time [8, 34]. Thus, simple, or parsimonious, explanations are preferred, which communicate the bulk of the model’s computation with a few concepts. However, inadequate faithfulness measures make the faithfulness-parsimony tradeoff hard to analyze, so prior U-CBEMs commonly set this hyperparameter arbitrarily (10 in ICE, 25 in CRAFT & ACE). SURF allows us to intelligently analyze this tradeoff.
Using the Object Classification task (Sec. 6.2), we fit U-CBEMs with an increasing number of concepts, each evaluated with the SURF measure and visualized in Fig. 2. Some U-CBEMs (CDISCO, CRAFT, C-SHAP, ICE) either marginally improve or perform worse on one or both metrics as they discover more concepts. SAEs initially improve but then oscillates. Only MCD and HU-MCD uniformly improve as the number of concepts increases, and they exhibit a plateauing effect, marking a natural choice for the number of concepts. Interestingly, HU-MCD is more faithful than MCD only when both discover few concepts; as the number of concepts grows, MCD becomes more faithful. This adds a nuanced touch to the findings from the HU-MCD paper, which claimed superior faithfulness over MCD based on results from a deletion-based proxy.
@@ 7 Conclusion
This paper argues that we lack a clear view on how faithful current U-CBEMs really are, largely because of measures that deviate from the formal definition of faithfulness. In accordance with this definition and associated desiderata, we propose SURF as a simple, principled measure for faithfulness. Among other things, we find that SOTA U-CBEMs do not faithfully explain the final output layer. While SURF accurately measures U-CBEM faithfulness, we emphasize that it should be paired with an interpretability analysis to judge the U-CBEM’s overall quality and value. We urge future work on concept-based explanations to adopt SURF as a standard faithfulness measure and to report SURF scores along their interpretability claims. U-CBEMs are predominantly applied on the final layer, but there is increasing interest in interpreting intermediate layers. SURF, as formulated, does not apply to this case, given the non-linear relationship between the explanation and the model’s output. Extending SURF to evaluate U-CBEMs for intermediate layers is important future work.
8 Acknowledgment
The support of the Office of Naval Research under grant N00014-24-1-2169, and IBM-Illinois Discovery Accelerator Institute (IIDAI), and USDA National Institute of Food and Agriculture under grant AFRI 2020-67021-32799/1024178 are gratefully acknowledged. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 21-46756. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Shubham Kumar also gratefully acknowledges the support of UIUC ECE’s Distinguished Research Fellowship and Promise of Excellence Fellowship.
@@ 8 Acknowledgment
    Sec. A: Discussion of related SAE faithfulness metrics.
@@
    Sec. B: Implementation details for the measure-over-measure study (Sec. 6.1).
  3. 3.
    Sec. C: Experimental details for the evaluation of U-CBEMs done in Sec. 6.2.
  4. 4.
    Sec. D: Results on additional models for the experiment done in Sec. 6.2.
A Discussion and Comparison with SAE Faithfulness Metrics
There has been a significant effort in recent years to explore using SAEs for explainability, particularly on large language models [5]. SAEs are typically trained to minimize embedding reconstruction error (e.g., mean squared error—MSE), which is the primary metric used to evaluate their faithfulness. The intuition is that if an SAE can accurately reconstruct the intermediate embedding it is explaining, then downstream operations on the reconstructed embedding should also be faithfully preserved. While reconstruction error is useful, it is a proxy for faithfulness; there is no clear mapping from reconstruction to downstream prediction error, making it hard to interpret the significance of a given MSE error. To accommodate this, recent work [16, 29] introduces downstream losses on model outputs to measure faithfulness. Common choices are KL divergence and cross entropy loss. However, we argue that these are uninterpretable (as acknowledged in [16]). These metrics work well when used relatively, which fits the referenced papers’s need to compare different SAEs, but these metrics do not give an absolute measure of faithfulness (unlike SURF), which is necessary when one wants to make an independent assessment of a method’s faithfulness.
Similarly, [31] develops a new SAE architecture (PatchSAE) and evaluates downstream classification accuracy after masking SAE latents (which is equivalent to a deletion-based approach). This metric suits their purpose, since they use SAEs to analyze model behavior, rather than to explain a prediction. [12] introduce a more robust SAE architecture (Archetypal SAEs) and introduce several metrics to evaluate. The closest metric to SURF is a plausibility metric, which measures the average maximum cosine similarity between the discovered CAVs and each classification vector. However, we argue that this metric does not give a holistic picture of faithfulness, since it only looks at the maximum and it is unclear how this metric maps on to errors in the output space (e.g., how does it get transformed by a softmax). We believe that SURF is a complementary measure that can be used to help study faithfulness of SAEs applied to vision models with the intention of explaining the model’s output computation.
B Implementation Details for Measure-over-Measure Comparison
@@
B.1 Experimental Setting Details
@@ B.2 Surrogate Implementation
B.2.1 SURF Surrogate (Ours)
@@ B.2.3 C-SHAP Surrogate
Since C-SHAP’s surrogate contains learnable parameters. Following [48], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities. We train the surrogate using the Adam optimizer [26] and a starting learning rate of $0.1$. The surrogate is trained for at most 100 epochs, during which the learning rate is decayed by a factor of 0.5 when the training loss plateaus. We stop training when the learning rate is reduced below $10^{-7}$.
B.3 FLOPs Computation
@@
C Additional Details for Benchmarking U-CBEM Faithfulness
@@ C.4 CDISCO
C.5 ICE
The open-source implementation provided by the authors was used to evaluate this approach. ICE uses NMF to find CAVs. Due to the potentially large number of embeddings, we use the batched-version of NMF. Concept importance is found using the TCAV method proposed in [24]. Embeddings are projected to the concept space using sci-kit learn’s transform method. ICE was extended to multi-attribute prediction by reformulating the task as independent binary classification tasks. Thus, CAVs and concept importances were learned for each task independently.
C.6 CRAFT
The open-source implementation provided in the Xplique toolbox [11] (by the same authors) was used to evaluate this approach. CRAFT uses NMF to find CAVs and introduces a Sobol-based sensitivity approach to find concept importance. Embeddings are projected to the concept space using the toolbox’s built-in transform method.
C.7 ConceptSHAP
@@
    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([33]) and is class-specific.
@@
Table 5: Additional Benchmark Results. We apply SURF to evaluate explanations from prior U-CBEMs on the Object Classification task from Sec. 6.2 on a VGG11 and InceptionV3. Along with the $\text{SURF}_{\text{MAE}}$ and $\text{SURF}_{\text{EMD}}$, we report other metrics to serve as a comparison. Our trends hold: prior U-CBEMs are not faithful, as indicated by large errors in the logit and probability space.
(a) Object classification
CDISCO 2.89 0.936 0.4 -0.01
ICE 3.58 0.616 93.7 0.498
CRAFT 2.50 0.844 51.0 0.214
C-SHAP 3.28 0.884 1.1 -0.00
MCD 1.95 0.738 83.3 0.306
HU-MCD
SAE 0.47 0.567 62.4 0.504
(b) Object classification
$\text{{SURF}}_{\text{MAE}}$ $\text{{SURF}}_{\text{EMD}}$ Top-1 (%) Rank Corr
4.08 0.847 1.52 0.012
3.34 0.595 50.3 -0.05
4.26 0.652 8.41 0.131
4.31 0.754 9.70 -0.08
2.76 0.327 89.1 0.385
0.35 0.226 72.0 0.770
C.8 MCD
Instead of representing concepts as single CAVs, MCD represents concepts as multi-dimensional subspaces. In [44], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace. To ensure the union of all concepts span the entire feature space, they define the orthogonal complement concept $C^{i,\perp}=\text{span}(C^{i,1},...,C^{i,K})^{\perp}$, where $K$ is the number of concepts originally discovered, letting $C^{i,K+1}\equiv C^{i,\perp}$. Then, any embedding $\mathbf{h}$ can be uniquely decomposed onto the concept basis as:
@@ C.9 HU-MCD
C.10 SAE
We learn a Top-K SAE [16] on the training embeddings (before average pooling) using the Overcomplete library. SAEs learn class-agnostic concepts (in contrast to U-CBEMs that learn CAVs for a specific class output); to enable a fair comparison, our SAE learns the same number of total CAVs as the other U-CBEMs for each setting. As an example, our SAE for the Object Classification task learns 505 CAVs (5 per class output, 101 total classes). At reconstruction time, we set the $K$ in the Top-K SAE to 5, directly mirroring the other U-CBEMs.
D Additional Benchmark Results
We repeat our experiment on the Object Classification task from Sec. 6.2 on a VGG11 and InceptionV3 to confirm trends hold. We use the ImageNet pre-trained weights, and finetune them for the Object Classification task. Results are reported in Tab. 5. For both models, SAE is the most faithful. For VGG11, we find that all U-CBEMs tend to exhibit large EMD errors (larger than 0.5). Though SAE is able to reduce the logit error significantly, the metrics indicate that there are still issues when trying to reproduce predictions in the probability space. While ICE can accurately reconstruct the Top-1 prediction, it has large logit errors. For InceptionV3, we find ICE is not faithful, and instead MCD is second-best (after SAE). Compared to the ResNet studied in Sec. 6.2, we believe that the models tested here have more nuance in their predicted probability space, which may lead to larger EMD errors, even if the logit error seems small.